# Entropic signature of resonant thermal transport: Ordered form of heat conduction

### v2

@@ Abstract
Thermal transport in crystals is influenced by chemistry, boundaries, and nanostructure. The phonon band structure extracted from molecular-dynamics simulations provides an illuminating view of both the type and extent of prevalence of wavelike mechanisms underlying the transport, yet falls short of elucidating the nature of thermal evolution for different phonon regimes. Here we present an analysis framework for the characterization of the entropic signature of the mechanisms induced by boundaries and nanostructure, using both equilibrium and nonequilibrium atomistic simulations. Specifically, we examine the effects of phonon confinement, Bragg scattering, and local resonances on the configurational phase space in room-temperature nanostructured silicon, and quantify how each modifies the rate of entropy production and thermal relaxation. We reveal that the presence of phonon local resonances spanning the full spectrum enables a highly ordered regime of heat conduction, where irreversible evolution and entropy maximization are severely hindered by extensive mode hybridizations caused by the resonances. This unique regime of transport paves the way for ultra-precise phonon control for a wide range of applications in condensed matter physics.
@@
In this Letter, we quantify--for the first time--the entropic signature of non-diffusive heat flow in the presence of dynamical constraints created by material nanostructuring, and we do so using MD simulations. The resulting framework bridges the phonon dispersion properties of any given class of nanostructures with the thermodynamic picture and offers a new route to distinguishing the coherent behavior of phonons. We consider a variety of crystalline nanoscale features that influence multiple phonon transport mechanisms at room temperature, including boundary scattering, confinement, Bragg scattering, and local resonances. Considering silicon as an example constituent material, for each system we first obtain the phonon band structure by calculating the spectral energy density of atomic motion obtained by MD simulation where anharmonic effects are accounted for. The dispersion relations highlight the influence of the type of nanostructure on the wavelike nature of phonons. Next we characterize the configurational entropy distribution at equilibrium and uncover non-homogeneous contributions influenced by the presence of the nanoscale features. We then combine nonequilibrium simulations and TUR to obtain a lower bound for the entropy production rate. This metric quantifies the impact of inherent dynamical phenomena on the production and exchange of entropy with the external environment, which allows us to characterize the strength of each configurational constraint in inhibiting the progression towards equilibrium. We find that NPMs are exceptional in this aspect owing to the unique ability of full-spectrum phonon-resonance hybridizations in hindering irreversible evolution. The existence of this kind of resonant phonon behavior, which manifests as a strongly fluctuating energy current, demonstrates that thermal energy currents do not necessarily correspond to entropy currents at the nanoscale. This brings to light a remarkable observation of reversibility in phonon transport in the presence of dynamical constraints.
Results and Discussion
For purposes of comparison, we consider the following geometries: bulk, a suspended membrane, a nanophononic crystal (NPC) with periodic internal cavities, and an NPM in the form of a nanopillared membrane (see geometric details in Fig. 1 and SM). We assume the Stillinger-Webber interatomic potential in all models [38]. Moreover, the simulated nanostructures are assumed to exhibit atomically flat surfaces. This idealization is made to isolate the aforementioned mechanisms and exclude the effects of nonpristine phonon-boundary scattering. Future work may examine the effects of surface roughness consistent with experimental conditions.
@@
In Fig. 1, we show the phonon spectrum for in-plane wave propagation obtained by applying SED analysis on equilibrium MD simulations [39, 14], as described in SM. Comparing the membrane, NPC, and NPM dispersions with the bulk case, we observe breakings of degeneracies and flattenings of the bands. In the NPM case, in particular, we observe profound changes to the phonon band structure due to the presence of a multitude of local resonances associated with the standing nanopillar [11, 40, 41, 14, 42, 13, 12, 15]. These resonances, also described as vibrons [42], appear in the form of horizontal branches at different frequencies spanning the full frequency range of the constituent material, in this case silicon. This feature represents the key character of an NPM. Its significance stems from the couplings with the underlying phonon dispersion curves associated with the base membrane, causing resonance hybridizations that reduce the group velocities of the heat carrying modes and provide a channel for phonon localization. These phonon band-structure features contribute to significant reduction in the in-plane thermal conductivity [42]. Also observed are some changes in the linewidths of the membrane, NPC, and NPM dispersion curves compared to the bulk case due to phonon-boundary scattering and lifetime reduction. While the SED spectrum reveals key information on the coherent and incoherent phonon dynamics, it fails to describe the nature of the thermodynamic evolution towards equilibrium associated with each nanostructure feature. This motivates the following investigation to formally characterize the entropic properties of each nanostructure.
@@
where $\rho$ is the local density around the atom, $k_{\mathrm{B}}$ is the Boltzmann constant, $g^{i}(r)$ is a smoothened radial distribution function, and $r_{m}$ is a cutoff radius [27]. For comparison purposes, $s^{i}$ can be interpreted as an estimation of the excess entropy of an atom relative to the ideal gas. Then, the total entropy of atom $i$ can be expressed as $S^{i}=s^{i}+S_{\rm IG}$, where the entropy per atom of the ideal gas $S_{\rm IG}$ is evaluated using the Sackur-Tetrode expression. Details on the evaluation of these expressions in equilibrium MD simulations are provided in SM, along with a convergence study and verification of the thermodynamic relation between $S$ and the internal energy.
In Fig. 2, we show local entropy maps for the different systems. Since we use the local density around the atom in Eq. ((1)), these calculations enable comparison of the available configurational phase space in different regions. We observe a larger accumulation of entropy close to the boundaries compared to the internal regions. Away from the boundaries, the crystalline order is enforced in all directions, thus strongly constraining each atom’s dynamics and reducing the size of the phase space of atomic configurations. In contrast, the atoms can accommodate a larger variety of microscopic configurations close to the boundaries due to the reduced number of neighbors. Consequently, the average equilibrium entropy $\bar{S}=\sum_{i}S^{i}/N$, where $N$ is the total number of atoms, increases with the surface-to-volume ratio, as clearly shown in the figure. This is consistent with the increased deviation of the dispersion from the bulk configuration. In addition, we observe that the entropy in the interior of the resonating nanopillars is significantly larger than in the interior of the membrane. We attribute this to the resonant modes in the nanopillar, which can correlate the dynamics of non-neighboring atoms, thus unlocking microstates that are not accessible in the membrane. Remarkably, this produces a significant increase in the size of the phase space in the NPM. We note that the observed differences in configurational entropy imply that each system needs to produce or exchange different amounts of entropy to reach equilibrium from given nonequilibrium conditions.
@@
Figure 3: Nonequilibrium entropy production (a) bound, and normalized (b) mean and (c) variance of the energy transferred between thermostats $q_{\tau}$ through the different material systems over a time-window $\tau$.
A microscopic configuration of a crystal lattice can be interpreted in the reciprocal representation in terms of the phonon population, which only maximizes the entropy when accommodating to the Bose-Einstein distribution. Therefore, the out-of-equilibrium phonon evolution in the different systems can be investigated in terms of entropy evolution. Here we use steady-state nonequilibrium MD simulations, where the terminals of the system are connected to Langevin thermostats at $T_{\text{C}}=290$ and $T_{\text{H}}=310$ K, respectively. To attenuate unphysical coupling/interaction between thermostats, we consider five unit cells between the thermal baths (see SM for details). In steady-state conditions, the entropy that the system produces internally is constantly exchanged with the thermostats, and quantifies the irreversible exchange of energy, or dissipation, to the system surroundings. According to the TUR [32], the thermodynamic uncertainty in terms of the energy current established in the system provides a lower bound on the entropy production $\Sigma_{\tau}$ during a time-window $\tau$:
@@
In such a stochastic interpretation of entropy production, the fraction of energy exchange with the thermostats that drive reversible phonon transitions in the system is not identified as energy dissipation and entropy exchange. Consistently, the entropy bound from the present MD simulations is smaller than the entropy exchange in the case of maximum energy dissipation (i.e., $\text{E}(q_{\tau})/T_{\text{C}}-\text{E}(q_{\tau})/T_{\text{H}}>\widehat{\Sigma}_{\tau}$) for all the cases considered, which in turn validates the TUR as a suitable tool in the present context [37, 47, 48]. Expression (3) is adequate to study the present systems far from equilibrium (where arbitrary non-Gaussian fluctuations may occur) beyond the linear-response interpretation [49]. Hence, $\widehat{\Sigma}_{\tau}$ is more robust than other metrics characterizing nanoscale nonequilibrium conditions such as the thermal conductivity, whose characterization requires strong assumptions. Nevertheless, it is illustrative to establish a relation between $\widehat{\Sigma}_{\tau}$ and an approximate bound for the thermal conductivity $k$ under the linear-response regime. Considering the entropy production rate in terms of the entropy fluxes $\Sigma_{\tau}/\tau=j/T_{\text{C}}-j/T_{\text{H}}$, with $j$ being the heat transfer rate, and assuming Fourier’s law, $\widehat{\Sigma}_{\tau}$ can be proportionally related to a lower bound on the conductivity [50], $\widehat{k}\eqsim(\widehat{\Sigma}_{\tau}T^{2})/(\tau V|\nabla{T}|^{2})$, where $V$ is the system volume, $T$ is the average temperature, and $\nabla T$ is the thermal gradient imposed between the thermostats.
@@
In Fig. 3(a), we plot the normalized entropy production bound $\widehat{\Sigma}_{\tau}/k_{\mathrm{B}}$ obtained from our nonequilibrium MD simulations considering $\tau=1$ ns. As demonstrated in SM, this is the smallest sampling time for which we obtain $\text{Var}(q_{\tau})/\text{E}({q}_{\tau})^{2}\sim 1/\tau$. This relation is a consequence of the central limit theorem in the long sampling time limit [51], and allows us to characterize a bound for the entropy production rate $\widehat{\Sigma}_{\tau}/\tau$ independent of $\tau$. Since $q_{\tau}$ fluctuates both in time and between different simulation realizations, we perform the averages during 30 ns over five different simulations in each case. Analysis of the computational size effects and the influence of the temperature difference between thermostats is provided in SM. As expected, the bulk is the system where entropy is produced most efficiently due to the absence of external constraints other than the thermostats. In contrast, the membrane exhibits a significant reduction in the entropy production rate. This is interpreted in terms of the local constraints to the phonon dynamics close to the free surfaces, which, in combination with the energy conservation restriction, inhibit some asymmetric transitions that are probable in the bulk case, thus reducing the overall probability of irreversible trajectories in the phase space that maximize the entropy. This is consistent with the lower $\widehat{\Sigma}_{\tau}$ observed in the NPC, where the addition of an internal cavity enhances the phonon-boundary interaction. In this case, however, we note that the entropy production reduction relative to bulk is too large to be solely attributed to more incoherent scattering due to an increase in the surface-to-volume ratio. Instead, the emergence of coherent interference of phonons, i.e., Bragg scattering, represents an additional constraint that can limit the efficient production of entropy. Finally, in the NPM, the addition of the nanopillars is shown to significantly reduce the production of entropy relative to the membrane even though no internal boundaries are included in the basis membrane, where the energy flows between the thermostats. Furthermore, the stronger the resonant effect due to a taller nanopillar, the lower the entropy the system can produce. Considering the relation between $\widehat{\Sigma}$ and $\widehat{k}$, this finding suggests that, despite the strong anharmonicity governing a silicon crystal, introducing local resonances across the full spectrum is an efficient strategy to manipulate nanoscale thermal transport [11, 52].
@@
The rate of entropy production also provides a means to quantitatively distinguish between the roles of coherent dynamical effects versus noncoherent effects. In Figs. 3(b) and 3(c), we show $\text{E}(q_{\tau})$ and $\text{Var}(q_{\tau})$, respectively. Noticeably, boundary scattering in the membrane and the NPC reduces the entropy production mainly by reducing the average energy flux, which also implies a strong variance reduction, while, in the NPM, $\widehat{\Sigma}_{\tau}$ is reduced mainly due to an increase in flux fluctuations. Specifically, we observe that the flux variance increases monotonically by enhancing the resonant effect, even though the average energy current is not significantly modified. According to Eqs. (2) and (3), these results imply that the energy transfer mechanisms in the presence of local resonances involve a reduced amount of irreversible internal transitions. The strong current fluctuations established in the system are thus signatures of the prevalence of coherent phonon effects unlocked by resonances  [^1] This is consistent with previous observation of an oscillating heat flux induced by other forms of resonance  [57]., while incoherent dynamics dominated by boundary scattering are characterized by a reduction in the energy transfer rate. Complementary analysis of this phenomena in terms of the flux autocorrelation at equilibrium is provided in SM.
Finally, we note that the ratio of the equilibrium entropy to the entropy production rate is a measure of the entropy maximization time. This metric characterizes the time required by a system to evolve from the minimum-entropy state, corresponding to a fully coherent and non-thermalized state, towards equilibrium at the given temperature. Even though quantification of this time would require integration of the entropy production through a specific nonequilibrium trajectory, we have shown that both $\Sigma_{\tau}$ and $\bar{S}$ are influenced by the resonant effects resident in NPMs even at room temperature. Therefore, the results indicate that the presence of local resonances delays the emergence of thermodynamic equilibrium and induce reversible phonon evolution. This unique finding offers a broader view that nanoscale thermal energy conduction in general is not necessarily a fully irreversible process. Leveraging these effects opens a new horizon for nanoscale phonon engineering with potential implications on a wide range of applications including semiconductor technology, thermoelectricity, superconductivity [54], and emerging quantum devices [55] that would benefit from controlled phonon evolution.

### v3

@@ Abstract
Thermal transport in crystals is influenced by chemistry, boundaries, and nanostructure. The anharmonic phonon band structure extracted from molecular-dynamics simulations provides an illuminating view of both the type and extent of prevalence of wavelike mechanisms underlying the transport, yet falls short of elucidating the nature of thermal evolution for different phonon regimes. Here we present an analysis framework for the characterization of the entropic signature of the mechanisms induced by boundaries and nanostructure, using both equilibrium and nonequilibrium atomistic simulations. Specifically, we examine the effects of phonon confinement, Bragg scattering, and local resonances on the configurational phase space in room-temperature nanostructured silicon, and quantify how each modifies the rate of entropy production and thermal relaxation. We reveal that the presence of phonon local resonances spanning the full spectrum enables a highly ordered regime of heat conduction, where irreversible evolution and entropy maximization are severely hindered by extensive mode hybridizations caused by the resonances. This unique regime of transport paves the way for ultra-precise phonon control for a wide range of applications in condensed matter physics.
I Introduction
@@
In this article, we quantify--for the first time--the entropic signature of non-diffusive heat flow in the presence of dynamical constraints created by material nanostructuring, and we do so using MD simulations. The resulting framework bridges the phonon dispersion properties of any given class of nanostructures with the thermodynamic picture and offers a new route to distinguishing the coherent behavior of phonons. We consider a variety of crystalline nanoscale features that influence multiple phonon transport mechanisms at room temperature, including boundary scattering, confinement, Bragg scattering, and local resonances. Considering silicon as an example constituent material, for each system we first obtain the phonon band structure by calculating the spectral energy density of atomic motion obtained by MD simulation where anharmonic effects are accounted for. The dispersion relations highlight the influence of the type of nanostructure on the wavelike nature of phonons. Next we characterize the configurational entropy distribution at equilibrium and uncover non-homogeneous contributions influenced by the presence of the nanoscale features. We then combine nonequilibrium simulations and TUR to obtain a lower bound for the entropy production rate. This metric quantifies the impact of inherent dynamical phenomena on the production and exchange of entropy with the external environment, which allows us to characterize the strength of each configurational constraint in inhibiting the progression towards equilibrium. We find that NPMs are exceptional in this aspect owing to the unique ability of full-spectrum phonon-resonance hybridizations in hindering irreversible evolution. The existence of this kind of resonant phonon behavior, which manifests as a strongly fluctuating energy current, demonstrates that thermal energy currents do not necessarily correspond to entropy currents at the nanoscale. This brings to light a remarkable observation of reversibility in phonon transport in the presence of dynamical constraints.
For purposes of comparison, we consider the following geometries: bulk, a suspended membrane, a nanophononic crystal (NPC) with periodic internal cavities, and an NPM in the form of a nanopillared membrane (see geometric details in Fig. 1 and Appendix A). We assume the Stillinger-Webber interatomic potential in all models [38]. Moreover, the simulated nanostructures are assumed to exhibit atomically flat surfaces. This idealization is made to isolate the aforementioned mechanisms and exclude the effects of nonpristine phonon-boundary scattering. Future work may examine the effects of surface roughness consistent with experimental conditions.
@@
II Anharmonic phonon dispersion relations
In Fig. 1, we show the phonon spectrum for in-plane wave propagation obtained by applying SED analysis on equilibrium MD simulations [39, 14], as described in Appendix B. Comparing the membrane, NPC, and NPM dispersions with the bulk case, we observe breakings of degeneracies and flattenings of the bands. In the NPM case, in particular, we observe profound changes to the phonon band structure due to the presence of a multitude of local resonances associated with the standing nanopillar [11, 40, 41, 14, 42, 13, 12, 15]. These resonances, also described as vibrons [42], appear in the form of horizontal branches at different frequencies spanning the full frequency range of the constituent material, in this case silicon. This feature represents the key character of an NPM. Its significance stems from the couplings with the underlying phonon dispersion curves associated with the base membrane, causing resonance hybridizations that reduce the group velocities of the heat carrying modes and provide a channel for phonon localization. These phonon band-structure features contribute to significant reduction in the in-plane thermal conductivity [42] and recently have been shown to cause distinct phonon localization along the direction of transport within anharmonic MD simulations [43]. Also observed are some changes in the linewidths of the membrane, NPC, and NPM dispersion curves compared to the bulk case due to phonon-boundary scattering and lifetime reduction. While the SED spectrum reveals key information on the coherent and incoherent phonon dynamics, it fails to describe the nature of the thermodynamic evolution towards equilibrium. This motivates the following investigation to formally characterize the entropic properties associated with each nanostructure feature.
@@ II Anharmonic phonon dispersion relations
III Equilibrium configurational entropy
@@
where $\rho$ is the local density around the atom, $k_{\mathrm{B}}$ is the Boltzmann constant, $g^{i}(r)$ is a smoothened radial distribution function, and $r_{m}$ is a cutoff radius [27]. For comparison purposes, $s^{i}$ can be interpreted as an estimation of the excess entropy of an atom relative to the ideal gas. Then, the total entropy of atom $i$ can be expressed as $S^{i}=s^{i}+S_{\rm IG}$, where the entropy per atom of the ideal gas $S_{\rm IG}$ is evaluated using the Sackur-Tetrode expression. Details on the evaluation of these expressions in equilibrium MD simulations at room temperature are provided in Appendix C, along with a convergence study and verification of the thermodynamic relation between $S$ and the internal energy.
In Fig. 2, we show local entropy maps for the different systems. Since we use the local density around the atom in Eq. (1), these calculations enable comparison of the available configurational phase space in different regions. We observe a larger accumulation of entropy close to the boundaries compared to the internal regions. Away from the boundaries, the crystalline order is enforced in all directions, thus strongly constraining each atom’s dynamics and reducing the size of the phase space of atomic configurations. In contrast, the atoms can accommodate a larger variety of microscopic configurations close to the boundaries due to the reduced number of neighbors. Consequently, the average equilibrium entropy $\bar{S}=\sum_{i}S^{i}/N$, where $N$ is the total number of atoms, increases with the surface-to-volume ratio, as clearly shown in the figure. This is consistent with the increased deviation of the dispersion from the bulk configuration. In addition, we observe that the entropy in the interior of the resonating nanopillars is significantly larger than in the interior of the membrane. We attribute this to the resonant modes in the nanopillar, which can correlate the dynamics of non-neighboring atoms, thus unlocking microstates that are not accessible in the membrane. Remarkably, this produces a significant increase in the size of the phase space in the NPM. We note that the observed differences in configurational entropy imply that each system needs to produce or exchange different amounts of entropy to reach equilibrium from given nonequilibrium conditions.
IV Entropy production
@@
A microscopic configuration of a crystal lattice can be interpreted in the reciprocal representation in terms of the phonon population, which only maximizes the entropy when accommodating to the Bose-Einstein distribution. Therefore, the out-of-equilibrium phonon evolution in the different systems can be investigated in terms of entropy evolution. Here we use steady-state nonequilibrium MD simulations, where the terminals of the system are connected to Langevin thermostats at $T_{\text{C}}=290$ and $T_{\text{H}}=310$ K, respectively. To attenuate unphysical coupling/interaction between thermostats, we consider five unit cells between the thermal baths (see Appendix D for details). In steady-state conditions, the entropy that the system produces internally is constantly exchanged with the thermostats, and quantifies the irreversible exchange of energy, or dissipation, to the system surroundings. According to the TUR [32], the thermodynamic uncertainty in terms of the energy current established in the system provides a lower bound on the entropy production $\Sigma_{\tau}$ during a time-window $\tau$:
@@
Figure 3: Nonequilibrium entropy production (a) bound, and normalized (b) mean and (c) variance of the energy transferred between thermostats $q_{\tau}$ through the different material systems over a time-window $\tau=$1 ns.
@@
In such a stochastic interpretation of entropy production, the fraction of energy exchange with the thermostats that drive reversible phonon transitions in the system is not identified as energy dissipation and entropy exchange. Consistently, the entropy bound from the present MD simulations is smaller than the entropy exchange in the case of maximum energy dissipation (i.e., $\text{E}(q_{\tau})/T_{\text{C}}-\text{E}(q_{\tau})/T_{\text{H}}>\widehat{\Sigma}_{\tau}$) for all the cases considered, which in turn validates the TUR as a suitable tool in the present context [37, 48, 49]. Expression (3) is adequate to study the present systems far from equilibrium (where arbitrary non-Gaussian fluctuations may occur) beyond the linear-response interpretation [50]. Hence, $\widehat{\Sigma}_{\tau}$ is more robust than other metrics characterizing nanoscale nonequilibrium conditions such as the thermal conductivity, whose characterization requires strong assumptions. Nevertheless, it is illustrative to establish a relation between $\widehat{\Sigma}_{\tau}$ and an approximate bound for the thermal conductivity $k$ under the linear-response regime. Considering the entropy production rate in terms of the entropy fluxes $\Sigma_{\tau}/\tau=j/T_{\text{C}}-j/T_{\text{H}}$, with $j$ being the heat transfer rate, and assuming Fourier’s law, $\widehat{\Sigma}_{\tau}$ can be proportionally related to a lower bound on the conductivity [51],
$\widehat{k}\eqsim\frac{\widehat{\Sigma}_{\tau}T^{2}}{\tau V|\nabla{T}|^{2}},$ (4)
where $V$ is the system volume, $T$ is the average temperature, and $\nabla T$ is the thermal gradient imposed between the thermostats.
@@
In Fig. 3(a), we plot the entropy production bound $\widehat{\Sigma}_{\tau}/k_{\mathrm{B}}$ obtained from our nonequilibrium MD simulations considering $\tau=1$ ns. As demonstrated in Appendix D, this is the smallest sampling time for which we obtain $\text{Var}(q_{\tau})/\text{E}({q}_{\tau})^{2}\sim 1/\tau$. This relation is a consequence of the central limit theorem in the long sampling time limit [52], and allows us to characterize a bound for the entropy production rate $\widehat{\Sigma}_{\tau}/\tau$ independent of $\tau$. Since $q_{\tau}$ fluctuates both in time and between different simulation realizations, we perform the averages during 30 ns over five different simulations in each case. Analysis of the computational size effects and the influence of the temperature difference between thermostats is provided in Appendix D. As expected, the bulk is the system where entropy is produced most efficiently due to the absence of external constraints other than the thermostats. In contrast, the membrane exhibits a significant reduction in the entropy production rate. This is interpreted in terms of the local constraints to the phonon dynamics close to the free surfaces, which, in combination with the energy conservation restriction, inhibit some asymmetric transitions that are probable in the bulk case, thus reducing the overall probability of irreversible trajectories in the phase space that maximize the entropy. This is consistent with the lower $\widehat{\Sigma}_{\tau}$ observed in the NPC, where the addition of an internal cavity enhances the phonon-boundary interaction. In this case, however, we note that the entropy production reduction relative to bulk is too large to be solely attributed to more incoherent scattering due to an increase in the surface-to-volume ratio. Instead, the emergence of coherent interference of phonons, i.e., Bragg scattering, represents an additional constraint that can limit the efficient production of entropy. Finally, in the NPM, the addition of the nanopillars is shown to significantly reduce the production of entropy relative to the membrane even though no internal boundaries are included in the basis membrane, where the energy flows between the thermostats. Furthermore, the stronger the resonant effect due to a taller nanopillar, the lower the entropy the system produces. Considering the relation between $\widehat{\Sigma}$ and $\widehat{k}$, this finding suggests that, despite the strong anharmonicity governing a silicon crystal, introducing local resonances across the full spectrum is an efficient strategy to manipulate nanoscale thermal transport [11, 53].
@@
The rate of entropy production also provides a means to quantitatively distinguish between the roles of coherent dynamical effects versus noncoherent effects. In Figs. 3(b) and 3(c), we show $\text{E}(q_{\tau})$ and $\text{Var}(q_{\tau})$, respectively. Noticeably, boundary scattering in the membrane and the NPC reduces the entropy production mainly by reducing the average energy flux, which also implies a strong variance reduction, while, in the NPM, $\widehat{\Sigma}_{\tau}$ is reduced mainly due to an increase in flux fluctuations. Specifically, we observe that the flux variance increases monotonically by enhancing the resonant effect, even though the average energy current is not significantly modified. The strong current fluctuations established in the system are thus signatures of the prevalence of coherent phonon effects unlocked by resonances, while incoherent dynamics dominated by boundary scattering are characterized by a reduction in the energy transfer rate. According to Eqs. (2) and (3), these results imply that the energy transfer mechanisms in the presence of local resonances involve a reduced amount of irreversible internal transitions. This unique finding offers a broader view that nanoscale thermal energy conduction in general is not necessarily a fully irreversible process.
V Heat flux autocorrelation in the presence of resonances
Figure 4: Heat flux ACF in NPMs. The heat flux is extracted from equilibrium simulations at room temperature for (a) the membrane and (b-e) the NPM with different nanopillar heights. We extract the heat flux from the entire simulation domains (b,d) and only from the base-membrane portion (c,e).
The strong current fluctuations in NPMs are associated with resonant hybridizations of the propagating phonon modes in the base membrane. In this section, we explore the origin and manifestation of this phenomenon using equilibrium MD simulations. Specifically, we first thermalize the NPM system following the protocol described in Appendix C. Thereafter and under the resulting equilibrium conditions at room temperature, we analyze the heat flux autocorrelation function (ACF) $<\textbf{J}(0)\otimes\textbf{J}(t)>$, where J is the heat flux vector integrated over the volume and sampled every 4 fs and $t$ is the delay time. For each nanostructure, we perform the simulations in the supercell consisting of five unit cells also considered in the previous section.
In Fig. 4, we show the heat flux ACF for different nanopillar heights $h$. First, we investigate the evolution of the global heat flux by integrating J over the volume of the entire NPM system. We observe that the autocorrelation of the thermal fluctuations decays completely over a time scale of 10s of picoseconds, which indicates that the atom dynamics in different regions of the system are not fully correlated. This suggests that thermodynamic equilibrium with ergodic evolution is robustly established in our simulations. Even though our simulation domains are relatively small and sustain strong dynamical constraints, the mixing condition is thus satisfied, which in turn justifies the applicability of the fluctuation theorem and the TUR.
To examine the influence of resonances in more detail, in Fig. 4 we also show the ACF by extracting the heat flux J only from the base-membrane portion of the NPM. The local heat flux ACF displays a remarkably slow initial decay followed by periodic oscillations that persist over a nanosecond time scale. This phenomenon is not observed in the bulk, membrane, or NPC systems, and becomes more pronounced in the presence of taller nanopillars, where the resonant effect is enhanced. Therefore, these features are a manifestation of coherent phonon dynamics in NPMs, and further confirm that the energy transport mechanisms in the base membrane are strongly influenced by the phonon coupling with the resonant modes in the nanopillars. In analogy with these results, previous work demonstrated that similar features in the heat flux ACF can appear in core-shell nanowires due to the emergence of resonant transversal modes that couple with the propagating longitudinal modes [54]--a mechanism that is distinctly different from the resonance hybridizations that occur in NPMs, yet still brings rise to this type of behavior.
Therefore, the anomalous behavior of the heat flux ACF under the influence of atomic-scale phonon resonances at equilibrium is a counterpart of the enhancement of the energy transfer fluctuations in non-equilibrium conditions (see Fig. 3). The coherent coupling of the phonon modes in the nanopillars and the base membrane impairs the stochastic evolution of the vibrational energy distribution, which hinders the emergence of thermodynamic equilibrium. This suggests that the ergodic exploration of the phase space is not as easily accessible in material systems presenting resonances compared to conventional materials, and might indicate a physical origin underlying errors and uncertainty in Green-Kubo calculations [55, 56] to extract the thermal conductivity in complex nanoscale configurations prone to various forms of resonances [54]. Nevertheless, quantification of the heat flux uncorrelation time and the conductivity was shown possible in NPMs with certain dimensions [57, 58] and other complex systems [59].
VI Entropy maximization time
The ratio of the equilibrium entropy obtained in Section III to the entropy production rate in Section IV is a measure of the entropy maximization time,
$\tau_{S}=\frac{N\bar{S}}{\widehat{\Sigma}_{\tau}/\tau}.$ (5)
This metric characterizes the time required by a system to evolve from the minimum-entropy state, corresponding to a fully coherent and non-thermalized state, towards equilibrium at the given temperature. However, $\tau_{S}$ does not correspond to actual thermalization times because the production rates $\widehat{\Sigma}_{\tau}/\tau$ are characterized in steady-state conditions at 300 K, whereas the entropy evolution is a function of temperature and sensitive to the specific non-equilibrium transient trajectory. Nevertheless, this idealized time enables equal-footing comparison of the influence of different dynamical constraints on the emergence of thermal equilibrium.
In Fig. 5, we show $\tau_{S}$ considering supercells consisting of five unit cells, and assuming a constant entropy production rate as characterized in Fig. 3. Accordingly, the modification of the equilibrium phase space and entropy production due to boundaries and nanostructure severely impacts the thermalization process. In particular, the results suggest that realization of NPM nanostructures exhibiting local phonon resonances is a route to profoundly hinder irreversible evolution and delay the emergence of equilibrium, which in turn opens the way for coherent, mechanistic, and precise manipulation of phonons.
Conclusion